Suppose you have an aunt who always eats 2 biscuits with one cup of tea. She eats biscuits only with tea. A can of biscuits costs $14 and it contains 20 biscuits. A cup of tea costs 60 cents. She reserves $60/month to her teas and biscuits. How many cups of tea will she drink per month?
Answer #1

20 biscuits would be eaten with 10 cups of tea.

Total expense = 14 + 0.6 x 10 = $ 20

Number of cups of tea in a month = 60/20 x 10 = 30 cups